# Auditing Human Decision-Making in High-Stakes Environments via Prescriptive AI: A Stress-Test on Real-Time Tactical Management

### v5

@@ Abstract
High-stakes decision-making is often compromised by cognitive biases and outcome dependency. Current AI models typically mimic historical human behavior, inheriting these biases and limiting their utility for normative improvement. Here, we introduce a Prescriptive AI framework designed to audit, rather than automate, human judgment in real-time environments. By decoupling decision quality from stochastic outcomes, we quantify ”decision latency” and status quo bias in elite soccer management—a high-pressure adversarial domain. Analyzing 2018 FIFA World Cup data, our system exposes critical risk states, such as performance collapse following salient positive events (e.g., an assist), which human experts systematically overlook due to outcome bias. These findings demonstrate that interpretable auditing systems can reveal structural flaws in human reasoning that predictive models obscure. This approach establishes a paradigm for Human-AI interaction prioritizing epistemic accountability over predictive mimicry in safety-critical domains.
@@ 1 Introduction
Despite these constraints, computational decision-support systems have been dominated by a fundamentally predictive paradigm. Models are typically optimized to replicate historical decision patterns, achieving statistical accuracy by mimicking prior human behavior. However, reliance on behavioral imitation is problematic: historical decisions often encode cognitive biases and institutional inertia. As a result, predictive systems create a structural mismatch between accuracy and decision quality, reaching a documented “predictive ceiling”—for instance, state-of-the-art substitution models plateau at approximately $70\%$ accuracy by merely cloning human choices (mohandas). Moreover, standard evaluation practices in stochastic environments suffer from outcome bias, judging decision quality by realized results rather than by the epistemic justification available at the time of choice (BaronHershey1988).
@@
Importantly, this work does not aim to improve predictive accuracy, automate decision-making, or replace human expertise. Instead, it focuses on auditing the quality of human judgment under uncertainty by separating epistemic justification from stochastic outcomes. From this perspective, the limitations of predictive systems are not merely due to data insufficiency, but reflect a normative flaw inherent to supervised learning in agentic contexts. By mapping contextual inputs to historical decisions, such systems inevitably reproduce embedded cognitive biases, including status quo bias (reluctance to change strategy) and sunk cost fallacy (persistence in failing courses of action). Consequently, predictive models tend to validate conservative human behavior rather than reveal when deviation is epistemically warranted.
@@
To address this limitation, we distinguish between Predictive AI, which forecasts events or replicates past decisions, and Prescriptive AI, defined here as a class of systems whose primary function is to audit, justify, and support human decision-making under uncertainty. This perspective aligns with decision-theoretic and epistemic accounts of agency, in which rational choice is defined as a coherent state transition at time $t$, independent of the stochastic outcome realized at time $t+n$ (vanBenthem2011). While prescriptive analytics and interpretable decision-support frameworks have been advocated in theory (Bertsimas2020; Lepenioti2020; Rudin2019), empirical demonstrations in realistic, adversarial environments remain scarce. Black-box models are particularly ill-suited for such auditing tasks due to automation bias and opacity (Parasuraman2010), motivating the need for systems that explicitly map observed states to epistemically justifiable actions.
@@
In this work, we use elite-level soccer as a demanding natural laboratory to operationalize this prescriptive auditing paradigm. Substitution decisions share structural properties with many safety-critical contexts: they are time-sensitive, irreversible, and made under pervasive uncertainty. Existing approaches in this domain are limited either by insufficient temporal resolution—introducing exposure bias by favoring cumulative playing time over efficiency (Pappalardo2019; SchmidtLilloBustos2024)—or by a reliance on purely predictive modeling that fails to capture normative tactical necessity (wu; pamukkale).
@@
To enable systematic auditing of human judgment, we propose a hybrid statistical–symbolic framework that overlays intrinsically interpretable fuzzy reasoning onto robust statistical signals. Rather than automating decisions, the system functions as an auditing layer, surfacing decision-relevant risk—such as performance decay or defensive liability—before it manifests as observable failure. Unlike descriptive fuzzy models (Marliere2017; zhou) or static ranking systems (salabun), the proposed approach operationalizes continuous, role-aware evaluation of decision states.
@@
This work makes three contributions. First, we formalize a prescriptive auditing framework for evaluating human decision-making under uncertainty, in which symbolic reasoning contextualizes statistical evidence rather than replacing human agency. Second, we introduce a role-aware cumulative mean metric that eliminates play-time exposure bias, enabling the detection of intra-match performance deterioration. Third, we provide empirical validation in a high-stakes adversarial setting, demonstrating that the proposed auditing system reveals latent risk patterns systematically overlooked by both human experts and black-box predictive models.
2 Results
2.1 Overcoming the Predictive Ceiling: Prescription vs. Prediction
Recent benchmarks in soccer analytics reveal a “predictive ceiling” for substitution modeling. mohandas analyzed 51,738 substitutions using supervised learning algorithms (Random Forest, SVM), achieving a maximum accuracy of $\approx 70\%$. This plateau suggests that nearly 30% of tactical decisions are driven by contextual nuances—such as fatigue thresholds or specific game plans—that remain opaque to black-box models.
Critically, supervised models in this domain suffer from a normative flaw: they are trained to mimic the human expert. If historical data encodes risk aversion or delayed reaction times, the ML model learns to replicate these biases. As detailed in Table 1, our proposed framework shifts the objective from minimizing prediction error to maximizing tactical utility, functioning as an auditor rather than an imitator.
Dimension Supervised ML Baseline (e.g., mohandas) Proposed Prescriptive System
2.2 The Mathematical Mechanism: Eliminating Exposure Bias
2.3 Stress-Testing Decision Latency: The Brazil vs. Belgium Case
2.3.1 Alignment and Early Detection
  • Willian (45’): Priority $P_{final}=72.0$, ranked 4th. Rule R04 (Rapid Drop) flagged a sharp decline in technical actions in the final 20 minutes of the first half. Immediate substitution by the coach confirms offensive ineffectiveness, supported by media reports ESPN_2018.
  • Gabriel Jesus (58’): Priority $P_{final}=99.1$, 2nd highest. Rule R08 (Ineffective Forward) triggered by low cumulative score and no goals/assists. Post-match data: 12 touches in 58 minutes Epoca_2018, validating tactical isolation.
  • Paulinho (73’): Priority $P_{final}=93.1$. System detected ”Early Fatigue” and low participation. External analysis noted minimal offensive impact Gazeta_2018.
2.3.2 The “Fagner Paradox”: Counterfactual Auditing
Table 2: Model Audit vs. Human Decision (Brazil vs. Belgium)
Player Analysis Slice($t$) $P_{\text{final}}$ Rank Diagnosis Real Decision
Willian 40’-45’ 72.0 4th Rapid Drop Substituted out
Gabriel Jesus 55’-60’ 99.1 2nd Ineffective Forward Substituted out
Fagner 45’-85’ 100.0 1st Critical Risk Not Substituted
Paulinho 65’-73’ 93.1 3rd Early Fatigue Substituted out
2.3.3 The “Lukaku Paradox”: Latency and Event Masking
Table 3: Comparative Analysis of Low-Volume Strikers
Player Context FCS Priority Human Response
G. Jesus (BRA) 12 Touches, 0 G/A High (#2) Substituted (58’)
R. Lukaku (BEL) 14 Touches, 1 Assist Max (#1) Delayed Sub (87’)
2.4 Systemic Validation: Longitudinal Patterns
Technical Anchors vs. Chronic Inefficacy.
Temporal Gap and Anticipation.
Post-Entry Validation.
2.5 Boundary Conditions: The Limits of Performance Auditing
3 Discussion
@@
  • Intrinsic interpretability as a structural requirement, not a post-hoc feature (Rudin2019);

### v6

# Prescriptive Artificial Intelligence: A Formal Paradigm for Auditing Human Decisions Under Uncertainty

@@ Abstract
We formalize Prescriptive Artificial Intelligence as a distinct paradigm for human–AI decision collaboration in high-stakes environments. Unlike predictive systems optimized for outcome accuracy, prescriptive systems are designed to recommend and audit human decisions under uncertainty, providing normative guidance while preserving human agency and accountability.
We introduce four domain-independent axioms characterizing prescriptive systems and prove fundamental separation results. Central among these is the Imitation Incompleteness theorem, which establishes that supervised learning from historical decisions cannot correct systematic decision biases in the absence of external normative signals. Consequently, performance in decision imitation is bounded by a structural bias term $\varepsilon_{\text{bias}}$ rather than the statistical learning rate $\mathcal{O}(1/\sqrt{n})$.
This result formalizes the empirically observed accuracy ceiling in human decision imitation tasks and provides a principled criterion for when automation should be replaced by epistemic auditing. We demonstrate the computational realizability of the framework through an interpretable fuzzy inference system, applied as a stress test in elite soccer decision-making, where it reveals systematic decision latency and risk states obscured by outcome and status quo biases.
The proposed framework establishes Prescriptive AI as a general, realizable class of decision-support systems applicable across safety-critical domains in which interpretability, contestability, and normative alignment are essential.
keywords:
Prescriptive Artificial Intelligence , Decision Auditing , Human–AI Collaboration , Epistemic State Transition , Outcome Bias , Fuzzy Inference Systems , Explainable Artificial Intelligence , Cognitive Bias , High-Stakes Decision Making , Normative Decision Theory
[…]
[inst1] organization=Institute of Computing, Universidade Federal Fluminense (UFF), addressline=Av. Gal. Milton Tavares de Souza, postcode=24210-346, city=Niterói, state=RJ, country=Brazil
@@ 1 Introduction
Despite these constraints, computational decision-support systems have been dominated by a fundamentally predictive paradigm. Models are typically optimized to replicate historical decision patterns, achieving statistical accuracy by mimicking prior human behavior. However, reliance on behavioral imitation is problematic: historical decisions often encode cognitive biases and institutional inertia. As a result, predictive systems create a structural mismatch between accuracy and decision quality, reaching a documented “predictive ceiling”—for instance, state-of-the-art substitution models plateau at approximately $70\%$ accuracy by merely cloning human choices [26]. Moreover, standard evaluation practices in stochastic environments suffer from outcome bias, judging decision quality by realized results rather than by the epistemic justification available at the time of choice [4].
@@
Importantly, this work does not aim to improve predictive accuracy, automate decision-making, or replace human expertise. Instead, it focuses on auditing the quality of human judgment under uncertainty by separating epistemic justification from stochastic outcomes. From this perspective, the limitations of predictive systems are not merely due to data insufficiency, but reflect a normative flaw inherent to supervised learning in agentic contexts. We argue that this limitation is fundamental: in the absence of external normative signals, decision imitation cannot, even asymptotically, correct systematic human bias. By mapping contextual inputs to historical decisions, such systems inevitably reproduce embedded cognitive biases, including status quo bias (reluctance to change strategy) and sunk cost fallacy (persistence in failing courses of action). Consequently, predictive models tend to validate conservative human behavior rather than reveal when deviation is epistemically warranted.
@@
To address this limitation, we distinguish between Predictive AI, which forecasts events or replicates past decisions, and Prescriptive AI, defined here as a class of systems whose primary function is to audit, justify, and support human decision-making under uncertainty. This perspective aligns with decision-theoretic and epistemic accounts of agency, in which rational choice is defined as a coherent state transition at time $t$, independent of the stochastic outcome realized at time $t+n$ [48]. While prescriptive analytics and interpretable decision-support frameworks have been advocated in theory [6, 21, 39], empirical demonstrations in realistic, adversarial environments remain scarce. Black-box models are particularly ill-suited for such auditing tasks due to automation bias and opacity [32], motivating the need for systems that explicitly map observed states to epistemically justifiable actions.
@@
In this work, we use elite-level soccer as a demanding natural laboratory to operationalize the prescriptive auditing paradigm. Substitution decisions share structural properties with many safety-critical contexts: they are time-sensitive, irreversible, and made under pervasive uncertainty. Existing approaches in this domain are limited either by insufficient temporal resolution—introducing exposure bias by favoring cumulative playing time over efficiency [30, 42]—or by a reliance on purely predictive modeling that fails to surface normative tactical risk [50, 28].
@@
To enable systematic auditing of human judgment, we propose a hybrid statistical–symbolic framework that overlays intrinsically interpretable fuzzy reasoning onto robust statistical signals. Rather than automating decisions, the system functions as an auditing layer, surfacing decision-relevant risk—such as performance decay or defensive liability—before it manifests as observable failure. Unlike descriptive fuzzy models [23, 52] or static ranking systems [40], the proposed approach operationalizes continuous, role-aware evaluation of evolving decision states.
@@
This work makes three contributions. First, we formalize a prescriptive auditing framework for evaluating human decision-making under uncertainty, grounded in domain-independent axioms and separation results that distinguish auditing from prediction. Second, we introduce a role-aware cumulative mean metric that eliminates play-time exposure bias, enabling principled detection of intra-episode performance deterioration. Third, we demonstrate the computational realizability of the proposed prescriptive framework in a high-stakes adversarial environment, showing that it systematically reveals latent risk patterns overlooked by both human experts and black-box predictive models.
2 Research Context and Review
This section situates the present work within the broader literature on decision theory, prescriptive analytics, and human-centered decision support. Beyond surveying domain-specific applications in sports analytics, the review deliberately incorporates foundational results from decision theory and cognitive science that formalize how decisions should be evaluated under uncertainty and how humans systematically deviate from normative rationality. These theoretical contributions provide the normative and cognitive grounding for interpreting Prescriptive AI not merely as an optimization technology, but as an auditing mechanism designed to compensate for structural limitations in human real-time decision-making.
2.1 Scope Summary of Reviewed Works
Category Representative Works
Theoretical Foundations
Prescriptive Analytics & Optimization [21]; [8]; [6]; [49]
Decision Support & Agency [36]; [29]; [44]; [4]; [32]
Logic of Agency & Information Dynamics [47]; [48]
Normative Decision Theory & Bounded Rationality [41]; [43]; [18]; [15]; [33]; [10]
Domain Applications
Contextual Tactical Reasoning [52]; [46]; [23]
Fuzzy Individual Evaluation [51]; [5]; [40]; [27]; [20]
Substitution Analysis & ML [16]; [37]; [28]; [50]; [26]
Performance Metrics [30]; [42]
Methodological Pillars
Explainable AI (XAI) [39]; [25]; [9]; [22]; [17]; [19]
Table 1: Scope-oriented categorization of related work, spanning theoretical foundations, domain-specific applications, and methodological pillars.
2.2 Theoretical Framework: Prescriptive AI and Decision Agency
To rigorously position the proposed system within the existing literature, it is necessary to distinguish between predictive and prescriptive paradigms of decision support. Prescriptive Analytics is commonly described as the highest stage of analytics maturity, addressing the question “what should be done?” through optimization, simulation, and rule-based reasoning [8, 21]. Recent systematic reviews confirm that while predictive models estimate future probabilities, prescriptive systems explicitly map observed states to recommended actions [49, 6], evolving historically from static Decision Support Systems (DSS) toward more dynamic and intelligent decision agents [29].
Despite this evolution, a critical gap persists in how prescriptive systems are evaluated in stochastic environments. Traditional approaches, including standard reinforcement learning (RL) agents [45], often exhibit outcome bias, in which the quality of a decision is judged primarily by its realized result rather than by the reasoning process that produced it [4]. In high-stakes domains involving human decision-makers, such retrospective evaluation is frequently insufficient. Moreover, the well-documented risk of automation bias, whereby users over-rely on opaque algorithmic recommendations [32], further motivates the need for prescriptive systems that function as interpretable decision auditors rather than black-box oracles.
At a conceptual level, these concerns resonate with work in the logical dynamics of information and agency, particularly van Benthem’s account of decision-making as an epistemic state transition under uncertainty [47, 48]. Within this tradition, rationality is not defined by the optimality of outcomes, but by the coherence of an action with respect to the informational state and constraints available at the moment the decision is taken. Decisions are thus evaluated ex ante, independently of the stochastic realization of downstream consequences.
This perspective provides a theoretical basis for separating decision quality at time $t$ from observed outcomes at time $t+n$, a distinction that has been increasingly explored across research communities concerned with decision agency, interpretability, and accountability. Rather than focusing exclusively on outcome prediction, these lines of work emphasize the importance of assessing whether decisions are contextually justified given the available evidence and risk structure, particularly in environments characterized by uncertainty, irreversibility, and delayed feedback.
From a normative standpoint, this ex-ante evaluation of decision quality is rooted in classical decision theory. Savage’s axiomatization of rational choice formalizes decision optimality as a function of expected utility conditional on the informational state available at time $t$, independently of future realizations [41]. From a normative standpoint, this ex-ante evaluation of decision quality is rooted in classical decision theory. Savage’s axiomatization of rational choice formalizes decision optimality as a function of expected utility conditional on the informational state available at time $t$, independently of future realizations [41]. In this sense, van Benthem defines the epistemic state underlying a decision, while Savage provides the normative evaluation over that state. Complementarily, the theory of bounded rationality introduced by Simon [43] demonstrates that human decision-makers are structurally incapable of performing such continuous optimization under time pressure. Empirical models of belief updating further show that sequential evidence integration is systematically biased by order and salience effects [18], while fast-and-frugal heuristics dominate real-time human judgment under uncertainty [15]. Together, these results reinforce the need for prescriptive systems that evaluate decision states normatively while preserving human agency, in line with established taxonomies of human–automation interaction [33] and theories of situation awareness [10].
2.3 Contextual Interpretation and Tactical Reasoning
Several studies model contextual indicators to support tactical interpretation in sports. [52] employ fuzzy contextual reasoning to produce action-oriented decisions, mapping continuous match descriptors into symbolic control outputs. Similarly, [23] applies fuzzy control systems to tactical arbitration under uncertainty. In contrast, [46] focus on the semantic recognition of tactical actions using fuzzy models, remaining descriptive rather than prescriptive. While these contributions advance contextual interpretation, they neither integrate individual performance evaluation nor address substitution decisions.
2.4 Fuzzy-Based Individual Performance Evaluation
A complementary line of research applies fuzzy logic to individual player evaluation. [51] and [5] employ fuzzy inference systems to assess player suitability and generate rankings. Subsequent works, such as [40] and [27], extend this paradigm through fuzzy and neuro-fuzzy models to synthesize performance-related variables. [20] further demonstrate the suitability of fuzzy logic for transforming physical indicators into interpretable assessments. These approaches provide interpretability but generally lack temporal dynamics or prescriptive substitution reasoning.
2.5 Substitution Analysis: From Prediction to Prescription
Research on player substitutions has traditionally followed two paths: observational analysis and predictive modeling. Studies such as [16], [37], [28], and [50] analyze substitutions using statistical, observational, or causal methods to understand timing and patterns. While valuable for post-match analysis, they do not directly support real-time decision-making.
More recently, [26] framed substitutions as a supervised prediction problem, applying machine learning models to estimate when a substitution is likely to occur. Despite achieving predictive accuracy of approximately 70%, this line of work emphasizes imitation of historical decisions rather than prescriptive optimization. Unlike Reinforcement Learning approaches that maximize a reward signal [45], which can be noisy in low-scoring sports, our approach focuses on the normative evaluation of the need for substitution, independent of the match’s stochastic outcome.
2.6 Performance Evaluation and Temporal Exposure Bias
A central challenge in multi-agent environments is quantifying individual contributions. The PlayeRank framework [30] defines a multidimensional metric validated against professional scouts. Although effective, its cumulative formulation reintroduces exposure bias, preventing the detection of performance decay or momentum reversals. Recent work by [42] demonstrates that player influence fluctuates meaningfully during a match due to tactical factors. Building on this, the present work adopts a role-aware cumulative mean over fixed temporal slices, addressing the exposure bias inherent in cumulative sums.
2.7 Symbolic Reasoning and Explainable AI (XAI)
The interpretation of imprecise information is critical for accountability. While “black-box” models dominate distinctive tasks, they are often unsuitable for high-stakes decision auditing [39]. Foundational works by [9], [22], and [25] emphasize the distinction between interpretability and post-hoc explanation. Surveys by [17] and [19] reinforce the need for transparency in human-centric AI. [39] specifically warns against explaining black boxes in high-stakes contexts, advocating for models where reasoning is transparent by design. This principle motivates our use of Fuzzy Logic not just as a controller, but as a semantic layer that ensures the system’s recommendations are contestable and auditible by the human coach.
2.8 Identified Research Gap
Despite substantial progress in both general analytics and sports science, the literature reveals three critical, intersecting gaps that this work addresses.
1. Lack of Formal Definition for Prescriptive AI: First and foremost, there is a lack of a formalized, normative definition for “Prescriptive AI.” While the term appears in recent industrial reports and applied solutions (e.g., [44]), the academic literature often conflates prescription with automation or simple prediction. There is virtually no framework that defines Prescriptive AI not just as a technological stack, but as a normative agent responsible for auditing the decision logic itself [21, 49].
2. Methodological Limitations in Dynamic Environments: Standard approaches to generating actionable insights, such as Reinforcement Learning, struggle with outcome bias [4]. In low-scoring, high-variance domains like soccer, a “good” decision can lead to a bad result (and vice-versa). Existing systems that rely on maximizing reward signals [45] or mimicking historical substitutions [26] fail to decouple the quality of the decision state from the stochastic outcome, rendering them unreliable for objective auditing.
3. Absence of Interpretable Decision Support in Soccer: Finally, the domain-specific literature remains fragmented. Soccer analytics focuses heavily on descriptive metrics [30] or predictive modelling, neglecting the prescriptive auditing of tactical decisions. No prior work jointly integrates (i) a temporally resolved, exposure-aware performance metric, (ii) intrinsic interpretability that mitigates automation bias [32], and (iii) a fuzzy reasoning engine capable of handling qualitative context.
The present work addresses these gaps by unifying temporal evaluation with symbolic decision modeling, effectively proposing a formal structure for Prescriptive AI: a system that evaluates decision quality via interpretable norms, independent of the stochastic match outcome.
3 A Normative Taxonomy of Action-Oriented Decision Systems: Formalizing Prescriptive AI
This section establishes a coherent conceptual hierarchy for action-oriented decision systems. We first characterize the broad and heterogeneous superset of Action-Oriented Decision Systems. We then situate Prescriptive Analytics as a structured decision-oriented framework within that space. Finally, we formally delineate Prescriptive Artificial Intelligence (Prescriptive AI) as a normative, high-stakes subset of Prescriptive Analytics, defined not by algorithmic choices but by its epistemological role in human decision-making.
Action-Oriented Decision Systems (The Heterogeneous Superset) Reinforcement Learning (Automation-Oriented) Recommender Systems Predictive Workflows Prescriptive Analytics (Decision-Oriented) Black-Box Optimization Simulation-Based Planning Prescriptive AIDecision AuditingIntrinsic InterpretabilityRisk-Aware Reasoning
Figure 1: A normative taxonomy of Action-Oriented Decision Systems The diagram illustrates the hierarchical relationship where Action-Oriented Decision Systems form the superset. Prescriptive Analytics sits within as a decision-focused subset, while Prescriptive AI serves as the normative core, defined by auditability and interpretability requirements.
3.0.1 Action-Oriented Decision Systems: A Heterogeneous Superset
  • 1.
    Optimization and mathematical programming systems;
  • 2.
    Simulation-based and what-if analysis tools;
  • 3.
    Rule-based and expert systems;
  • 4.
    Reinforcement Learning (RL) agents;
  • 5.
    Recommendation and next-best-action systems;
  • 6.
3.0.2 Prescriptive Analytics as a Decision-Oriented Subset
Within this heterogeneous landscape, Prescriptive Analytics provides a more structured and widely recognized framework. Originating in business intelligence and operations research, prescriptive analytics focuses on determining what actions should be taken to achieve specified objectives, typically through optimization, simulation, and rule-based reasoning [8, 21, 6].
Prescriptive Analytics narrows the scope of Action-Oriented Decision Systems by explicitly addressing decision recommendation. However, it remains largely agnostic with respect to how uncertainty, human judgment, or accountability should be handled. It encompasses both automated and decision-support systems and does not inherently distinguish between optimizing outcomes and supporting human deliberation.
Accordingly, Prescriptive Analytics is best understood as a decision-oriented subset of Action-Oriented Decision Systems: analytically grounded, but normatively under-specified.
3.0.3 Formal Definition of Prescriptive AI
Definition
Prescriptive AI is a subclass of what is traditionally referred to as prescriptive analytics within the broader class of Action-Oriented Decision Systems, formally characterized by a normative decision operator whose role is to determine optimal actions and audit decision coherence under uncertainty, rather than to automate decisions or predict outcomes. This operator admits two equivalent temporal instantiations.
Norm-Invariant Recommendation and Auditing
Prescriptive AI employs a single normative decision framework for both recommendation and auditing. In prospective use, the system operates online, producing action recommendations from the epistemic state available at decision time. In retrospective analysis, the same framework is applied offline to historical records in a backtesting setting.
Auditing consists of reconstructing, from logged data, the information that was available at the moment a decision was made and recomputing the recommendation that would have been generated under the same normative criteria. Future observations are used exclusively to denoise or complete past measurements (e.g., correcting sensor errors or filling missing values), and never to introduce outcome knowledge.
Decision coherence is then assessed by comparing the reconstructed normative recommendation with the action actually taken. No separate temporal operator or learning procedure is introduced: auditing is simply norm-invariant backtesting. The same prescriptive logic governs both real-time recommendation and retrospective evaluation; only the data access mode differs.
Prescriptive AI systems are not characterized by the techniques they employ, but by the normative role they play in the decision process. In particular, they are distinguished by the following defining properties:
  • 1.
    A primary function of decision recommendation or decision recommendation rather than decision automation;
  • 2.
@@ 3.0.3 Formal Definition of Prescriptive AI
  • 3.
    Intrinsic interpretability as a structural requirement, not a post-hoc feature [39];
  • 4.